Ellis Transport Company acquired 1.7 million shares of stock in L&K Corporation at $77 per share. They are classified by Ellis as “available for sale.” Ellis sold 700,000 shares at $80, received a 10% stock dividend, and then later in the year sold another 200,000 shares at $76.

Hint: There is no entry for the stock dividend, but a new investment per share must be calculated for use later when the shares are sold.

Required:
Prepare journal entries to record these transactions. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field. Enter your answers in millions rounded to 1 decimal place (i.e., 5,500,000 should be entered as 5.5).)
Answer #1

Answer:

Date Account Debit credit
1 Investment in availabe for sale securities 130900000
cash 130900000
[being investment made 1,400,000*88]
2 cash [700000*80] 56000000
Gain on sale of available for sale security 2100000
Investment in available for sale security [700000*77] 53900000
[Being investment sold at a gain]
3 cash   [200000*76] 15200000
Gain on sale of available for sale security 2400000
Investment in available for sale security [200000*64] 12800000
[being investment sold at gain]

Number of shares outstanding after first sale = 1,700,000-700,000 = 1,000,000 shares

Stock dividend received : 1,000,000*.10= 100,000 shares

Total shares outstanding : 1000000+200000= 1,200,000 shares

Value of investment after first sale : [130,900,000 - 53,900,000]

          = 77,000,000/1,200,000= $ 64 per share

**the stock dividend received reduces the cost of investment